# UniMODE: Unified Monocular 3D Object Detection

### v4

@@ 4.1 Performance Comparison
According to the results, we can observe that UniMODE achieves the best results in all metrics. It surpasses the SOTA Cube RCNN by 4.9% given the primary metric $\rm AP_{3D}$. Besides DLA34, we also try another backbone ConvNext-Base. This is because previous papers suggest that DLA34 is commonly used in camera-view detectors like Cube RCNN but unsuitable for BEV detectors [13]. Since UniMODE is a BEV detector, testing its performance with only DLA34 is unfair. Thus, we also test UniMODE with ConvNext-Base, and the result suggests that the performance is boosted significantly. Additionally, the speed of UniMODE is also promising. Test on 1 A100 GPU, the inference speeds of UniMODE under the high and low computing resource settings are 21.41 FPS and 43.48 FPS separately. Moreover, it can be observed from Table 1 that BEVFormer and PETR do not converge well in unified detection while behaving promisingly trained with outdoor datasets. This phenomenon implies the difficulty of unifying indoor and outdoor 3D object detection. Through analysis, we find that BEVFormer obtains poor results when using data of all domains because its convergence is quite unstable, and the loss curve often boosts to a high value during training. PETR does not behave well since it implicitly learns the correspondence relation between 2D pixels and 3D voxels. When the camera parameters keep similar across all samples in a dataset like nuScenes [6], PETR converges smoothly. Nevertheless, when trained in a dataset with dramatically changing camera parameters like Omni3D, PETR becomes much more difficult to train.
4.2 Ablation Studies
@@
Table 3: Ablation studies on proposed strategies, which verify the effects of proposal head (PH), uneven BEV grid (UBG), sparse BEV feature projection(SBFP), and unified domain alignment (UDA). The last column presents the improvement of each row compared with the top row. $\rm AP_{3D}^{in}$ and $\rm AP_{3D}^{out}$ reflect the indoor and outdoor detection performance, respectively. Notably, although SBFP does not boost detection precision, it reduces the computational cost of the BEV feature projection by 82.6%.
Key component designs. We ablate the effectiveness of the proposed strategies, including proposal head, uneven BEV grid, sparse BEV feature projection, and unified domain alignment, and present the results in Table 3.
According to the results in Table 3, we can observe that all these strategies are very effective. Among them, the proposal head boosts the result by the most significant margin. Specifically, the proposal head enhances the overall detection performance metric $\rm AP_{3D}$ by 3.6%. Meanwhile, the indoor and outdoor detection metrics $\rm AP_{3D}^{in}$ and $\rm AP_{3D}^{out}$ are boosted by 2.5% and 7.9% separately. As discussed in Section 3.2, the proposal head is quite effective because it stabilizes the convergence process of UniMODE and thus favors detection accuracy. The collapse does not happen after using the proposal head. In addition, although the sparse BEV feature projection strategy does not improve the detection precision, it reduces the projection cost by 82.6%.
Figure 5: Visualization of detection results on various sub-datasets in Omni3D.
@@
Table 4: Ablation studies on grid size and depth bin split strategy in uneven BEV grid.
@@
Table 5: Ablation study on $\tau$ in sparse BEV feature projection.
Uneven BEV grid. We study the effect of BEV feature grid size and depth bin split strategy in uneven BEV grid design, and the results are presented in Table 4. When the depth bin split is uneven, we split the depth bin range following Eq. 1. Comparing the $1_{\rm st}$ and $2_{\rm nd}$ rows of results in Table 4, we can find that uneven depth bin deteriorates detection performance. We speculate this is because this strategy projects more points in closer BEV grids while fewer points in farther grids, which further increases the imbalanced distribution of projection features. Additionally, comparing the $1_{\rm st}$, $3_{\rm rd}$, and $4_{\rm th}$ rows of results in Table 4, it is observed that smaller BEV grids lead to better performance.
Sparse BEV feature projection. As mentioned in Section 3.4, the BEV feature projection process is computationally expensive. To reduce this cost, we propose to remove unimportant projection points. Although this strategy enhances network efficiency significantly, it could deteriorate detection accuracy and convergence stability, which is exactly a trade-off. In this part, we study this trade-off through experiments. Specifically, as introduced in Section 3.4, we remove unimportant projection points based on a pre-defined hyper-parameter $\tau$. The value of $\tau$ is adjusted to analyze how the removed projection point ratio affects performance. The results are reported in Table 5.
It can be observed from Table 5 that when $\tau$ is 0, which means no feature is discarded, the best performance across all rows is arrived. When we set $\tau$ to $1e^{-3}$, about 82.6% of the feature is discarded while the performance of the detector remains very similar to the one with $\tau=0$. This phenomenon suggests that the discarded feature is unimportant for final detection accuracy. Then, when we increase $\tau$ to $1e^{-2}$ and $1e^{-1}$, we can find that the corresponding performances drop dramatically. This observation indicates that when we discard superfluous features, the detection precision and even training stability are influenced significantly. Combining all the observations, we set $\tau$ to $1e^{-3}$ and drop $82.6\%$ of unimportant features in UniMODE, which reduces the computational cost by $82.6\%$ while maintaining similar performance as the one without dropping features.
@@
Table 6: Analysis on the effectiveness of DALN.
Effectiveness of DALN. In this experiment, we validate the effectiveness of DALN through comparing the performances of the naive baseline without any domain adaptive strategy, the baseline predicting dynamic parameters with direct regression (DR) [24], and the baseline with DALN (our proposed). All these models are trained using only ARKitScenes and evaluated with ARKitScenes (in-domain) and SUN-RGBD (out-of-domain) separately. The result is presented in Table 6. It can be observed that DR could degrade the detection accuracy while DALN boosts the performance significantly, which reveals the zero-shot out-of-domain effectiveness of DALN.
4.3 Cross-domain Evaluation
We evaluate the generalization ability of UniMODE in this part by conducting the cross-domain evaluation. Specifically, we train a detector on a sub-dataset in Omni3D and test the performance of this detector on different other sub-datasets. The experiments are conducted in two settings. In the zero-shot setting, the test domain is completely unseen. In the $\delta$-tune setting, 1% of training set data from the test domain is used to fine-tune the Query FFN in UniMODE for 1 epoch. The experimental results are presented in Table 7.
@@ 4.3 Cross-domain Evaluation
Table 7: Cross-domain evaluation in indoor sub-datasets. In this experiment, the detector is first trained in a domain and tested on other domains in two settings, zero-shot and $\delta$-tune.
@@
According to the results in the $2_{\rm st}\sim 4_{\rm th}$ columns of Table 7, we can find that when a detector is trained and validated in the same indoor sub-dataset, its performance is promising. However, when evaluated on another completely unseen sub-dataset, the accuracy is limited, especially for the virtual dataset Hypersim. This is partly because monocular 3D depth estimation is an ill-posed problem. The results of the $\delta$-tuning setting are reported in the $5_{\rm st}\sim 7_{\rm th}$ columns of Table 7. We can observe that when fine-tuned with only a handful of data, the performance of UniMODE becomes much more promising.
4.4 Visualization
We visualize the detection results of UniMODE on various sub-datasets in Omni3D. The illustrated results are shown in Fig. 5, where UniMODE performs well on all the data samples, and accurately captures the 3D object bounding boxes under both complex indoor and outdoor scenarios.
5 Conclusion and Limitation
In this work, we have proposed a unified monocular 3D object detector named UniMODE, which contains several well-designed techniques to address many challenges observed in unified 3D object detection. The proposed detector has achieved SOTA performance on the Omni3D benchmark and presented high efficiency. The limitation of the detector is its zero-shot generalization ability on unseen data scenarios is still limited.
Acknowledgements. This work is supported by National Natural Science Foundation of China (No. 62201484), Meta Open Science Research Fund, HKU Startup Fund, HKU Seed Fund for Basic Research, and Natural Science Foundation of Zhejiang Province (No. LD24F020002).

### v5

@@ V-A Monocular Performance Comparison
According to the results, we can observe that UniMODE achieves the best results in all metrics. It surpasses the SOTA Cube RCNN by 4.9% given the primary metric $\rm AP_{3D}$. Besides DLA34, we also try another backbone ConvNext-Base. This is because previous papers suggest that DLA34 is commonly used in camera-view detectors like Cube RCNN but unsuitable for BEV detectors [52]. Since UniMODE is a BEV detector, testing its performance with only DLA34 is unfair. Thus, we also test UniMODE with ConvNext-Base, and the result suggests that the performance is boosted significantly. According to the results in the last two rows in Table II, the detector with ConvNext-base outperforms the detector with DLA34 by 2.7% $\rm AP_{3D}$. Additionally, the speed of UniMODE is also promising. Test on 1 A100 GPU, the inference speeds of UniMODE under the high and low computing resource settings are 21.41 FPS and 43.48 FPS separately. Moreover, it can be observed from Table I that BEVFormer and PETR do not converge well in unified detection while behaving promisingly trained with outdoor datasets. This phenomenon implies the difficulty of unifying indoor and outdoor 3D object detection. Through analysis, we find that BEVFormer obtains poor results when using data of all domains because its convergence is quite unstable, and the loss curve often boosts to a high value during training. PETR does not behave well since it implicitly learns the correspondence relation between 2D pixels and 3D voxels. When the camera parameters keep similar across all samples in a dataset like nuScenes [43], PETR converges smoothly. Nevertheless, when trained in a dataset with dramatically changing camera parameters like Omni3D, PETR becomes much more difficult to train.
V-B Multi-modal Performance Comparison
In this experiment, we compare the performance MM-UniMODE with previous multi-modal counterparts on the MM-Omni3D benchmark. Since there do not exist detectors that are designed for the unified multi-modal 3D object detection setting in MM-Omni3D, we select three representative methods (including CMT [30], BEVFusion [29], and FusionFormer [50]) and reproduce them in MM-Omni3D. Among them, CMT adopts a detection architecture similar to the original DETR, while BEVFusion and FusionFormer are convolution-based BEV methods. All the detectors are validated with two backbones, ConvNext-Base and VoVNet [53]. The results are presented in Table III. The result of UniMODE is also presented for the convenience of comparison. As mentioned before, the experiments in Table III are performed in the low computing resource setting. The best result of every metric is marked in bold. We can find that MM-UniMODE obtains the best performance in all metrics. The results of MM-UniMODE are significantly better than UniMODE, which confirms that incorporating depth sensor information as input benefits detection accuracy. Additionally, although compared with ConvNext-Base, the VoVNet backbone improves the outdoor results of BEVFusion, FusionFormer, and MM-UniMODE significantly, VoVNet degrades the precision of CMT, which is mainly because of the model structure difference. Moreover, comparing the results in Table I and Table III, although CMT is a multi-modal detector extended from the monocular detector PETR, the training of PETR collapses while CMT achieves much better results, which implies that the depth information can stabilize the training of unified 3D object detectors. We speculate this is because the depth information addresses the ill-posed nature of monocular 3D object detection.
TABLE IV: Ablation studies on the proposed strategies in UniMODE.
@@ V-B Multi-modal Performance Comparison
V-C Ablation Studies
UniMODE key designs. In this experiment, we ablate the effectiveness of the proposed strategies in UniMODE, including proposal head (PH), uneven BEV grid (UBG), sparse BEV feature projection (SBFP), and unified domain alignment (UDA). The experimental results are reported in Table IV. Specifically, the last column of Table IV presents the improvement of each row compared with the top row. $\rm AP_{3D}^{in}$ and $\rm AP_{3D}^{out}$ reflect the indoor and outdoor detection performance, respectively. Notably, although SBFP does not boost detection precision, it reduces the computational cost of the BEV feature projection by 82.6%. Overall, we can observe that all these strategies are very effective. Among them, the proposal head boosts the result by the most significant margin. Specifically, the proposal head enhances the overall detection performance metric $\rm AP_{3D}$ by 3.6%. Meanwhile, the indoor and outdoor detection metrics $\rm AP_{3D}^{in}$ and $\rm AP_{3D}^{out}$ are boosted by 2.5% and 7.9% separately. As discussed in Section III-B, the proposal head is quite effective because it stabilizes the convergence process of UniMODE and thus favors detection accuracy. The collapse does not happen after using the proposal head. In addition, UBG and UDA are two plug-and-play strategies, which are not only effective in improving performance but also easy to implement. Furthermore, although the sparse BEV feature projection strategy does not improve the detection precision, it reduces the projection cost by 82.6%.
TABLE V: Ablation studies on grid size and depth bin split strategy.
@@ V-C Ablation Studies
TABLE VI: Ablation study on $\tau$ in sparse BEV feature projection.
@@
Analysis of uneven BEV grid. We study the effect of BEV feature grid size and depth bin split strategy in uneven BEV grid design, and the results are presented in Table V. When the depth bin split is uneven, we split the depth bin range following Eq. 1. Comparing the $1_{\rm st}$ and $2_{\rm nd}$ rows of results in Table V, we can find that uneven depth bin deteriorates detection performance. We speculate this is because this strategy projects more points in closer BEV grids while fewer points in farther grids, which further increases the imbalanced distribution of projection features. Additionally, comparing the $1_{\rm st}$, $3_{\rm rd}$, and $4_{\rm th}$ rows of results in Table V, it is observed that smaller BEV grids lead to better performance. Due to the consideration of computing resource budget, We set the BEV grid size to 1m rather than 0.5m in all the other experiments. According to the results of this experiment, we can infer that if we further decrease the size of BEV grids, the performance of UniMODE can be continuously boosted compared with the current performance.
Analysis of sparse BEV feature projection. As mentioned in Section III-D, the BEV feature projection process is computationally expensive. To reduce this cost, we propose to remove unimportant projection points. Although this strategy enhances network efficiency significantly, it could deteriorate detection accuracy and convergence stability due to the reduction of retained feature. Hence, this is exactly a trade-off between efficiency and detection precision. In this part, we study this trade-off through experiments. Specifically, as introduced in Section III-D, we remove unimportant projection points based on a pre-defined hyper-parameter $\tau$. Therefore, in this experiment, we adjust the value of $\tau$ to analyze how the removed projection point ratio affects performance, and the results are reported in Table VI.
It can be observed from Table VI that when $\tau$ is 0, which means no feature is discarded, the best performance across all rows is arrived. When we set $\tau$ to $1e^{-3}$, about 82.6% of the feature is discarded while the performance of the detector remains very similar to the one with $\tau=0$. This phenomenon suggests that the discarded feature in the setting $\tau=1e^{-3}$ is unimportant for maintaining the final detection accuracy. Then, when we increase $\tau$ to $1e^{-2}$ and $1e^{-1}$, we can find that the corresponding detection performances drop dramatically. This observation indicates that when we discard superfluous features, the detection precision and even training stability would be influenced significantly. Combining all the observations, we set $\tau$ to $1e^{-3}$ and drop $82.6\%$ of unimportant features in UniMODE, which reduces the computational cost by $82.6\%$ while maintaining similar performance as the one without dropping features.
TABLE VII: Analysis on the effectiveness of DALN.
@@
TABLE VIII: Ablation study on the MIC strategy.
$L_{P\rightarrow I}$ $L_{I\rightarrow P}$ ${\rm AP_{3D}^{sun}}$ ${\rm AP_{3D}^{hyp}}$ ${\rm AP_{3D}^{ark}}$ ${\rm AP_{3D}^{obj}}$ ${\rm AP_{3D}^{kit}}$ ${\rm AP_{3D}^{nus}}$ ${\rm AP_{3D}}$
16.5% 5.7% 36.7% 50.8% 20.5% 23.8% 23.8%
16.3% 5.9% 37.8% 56.0% 22.7% 27.4% 25.1%
17.1% 5.9% 38.3% 56.4% 22.7% 26.3% 25.7%
17.1% 6.1% 39.2% 57.1% 23.1% 26.2% 26.1%
Analysis of DALN. In this experiment, we validate the effectiveness of DALN through comparing the performances of the naive baseline without any domain adaptive strategy, the baseline predicting dynamic parameters with direct regression (DR) [54], and the baseline with DALN (our proposed). All these models are trained using only the ARKitScenes dataset. Afterwards, they are evaluated on ARKitScenes (in-domain evaluation) and SUN-RGBD (out-of-domain evaluation) separately. The experimental result is presented in Table VII. It can be observed that DR could degrade the detection accuracy. This is because directly predicting accurate dynamic parameters is challenging and harms the training stability. By contrast, DALN boosts the performance significantly, which reveals the zero-shot out-of-domain effectiveness of DALN.
Analysis of MIC. In this part, we verify the effectiveness of MIC used for image and point feature fusion in MM-UniMODE. According to the description in Section IV-D, there are two kinds of guidance losses in MIC, i.e., $L_{P\rightarrow I}$ and $L_{I\rightarrow P}$. Among them, $L_{P\rightarrow I}$ utilizes precise depth information to guide the extraction of ambiguous image feature, and $L_{I\rightarrow P}$ employs informative image representation to aid the feature lack in grids that do not contain depth sensor points. In this experiment, we evaluate the effectiveness of $L_{P\rightarrow I}$ and $L_{I\rightarrow P}$ individually. The results are reported in Table VIII. As presented, both $L_{P\rightarrow I}$ and $L_{I\rightarrow P}$ boost the baseline detector performance significantly. Among them, $L_{I\rightarrow P}$ plays a more important role, which boosts the ${\rm AP_{3D}}$ metric by 1.9%. Moreover, $L_{P\rightarrow I}$ enhances ${\rm AP_{3D}}$ by 1.3%, which is also a significant improvement. When we combine $L_{I\rightarrow P}$ and $L_{P\rightarrow I}$, the overall detection precision is improved by 2.3%. The results well confirm the effectiveness of the proposed MIC module and guidance loss items.
TABLE IX: Cross-domain evaluation in indoor sub-datasets.
@@
TABLE X: Analysis on how training data volume affects out-of-domain generalization performance.
Training Data ${\rm AP_{3D}^{sun}}\uparrow$
SUN-RGBD 17.3%
SUN-RGBD & Objectron & ARKitScenes 21.5%
Objectron & ARKitScenes 19.6%
Objectron & ARKitScenes & KITTI 20.1%
TABLE XI: Analysis on how MM-Omni3D pre-training affects the performance in the KITTI test set.
Training ${\rm AP_{3D}^{easy}}\uparrow$ ${\rm AP_{3D}^{mode.}}\uparrow$ ${\rm AP_{3D}^{hard}}\uparrow$
KITTI 90.0% 81.2% 75.4%
DD3D-KITTI 89.8% 82.6% 79.4%
MM-Omni3D 91.2% 84.8% 80.7%
MM-Omni3D* 91.9% 85.2% 81.2%
V-D Monocular Cross-domain Evaluation
We evaluate the generalization ability of UniMODE in this part by conducting cross-domain evaluation between various indoor sub-datasets in Omni3D. Specifically, we train a detector on an indoor sub-dataset of Omni3D and test the performance of this detector on different other sub-datasets. The experiments are conducted in two settings, i.e., the zero-shot setting and $\delta$-tune setting. In the zero-shot setting, the test domain is completely unseen. In the $\delta$-tune setting, 1% of the training set data from the test domain is used to fine-tune the Query FFN in UniMODE for 1 epoch. The experimental results are presented in Table IX.
@@ V-D Monocular Cross-domain Evaluation
According to the results in the $2_{\rm st}\sim 4_{\rm th}$ columns of Table IX, we can find that when a detector is trained and validated in the same indoor sub-dataset, its performance is promising. However, when evaluated on another completely unseen sub-dataset, the accuracy is limited, especially for the virtual dataset Hypersim. This is partly because monocular 3D depth estimation is an ill-posed problem. When the training and validation data belong to differing domains, predicting depth accurately is challenging, especially for the virtual dataset Hypersim. Then, we introduce another testing setting, $\delta$-tuning. In this setting, if a detector is tested on another domain different from the training domain, the Query FFN is fine-tuned by 1% of training data from the test domain. The results of the $\delta$-tuning setting are reported in the $5_{\rm st}\sim 7_{\rm th}$ columns of Table IX. We can observe that when fine-tuned with only a handful of data, the performance of UniMODE becomes much more promising. This result suggests that the superiority of UniMODE in serving as a foundational model. It can benefit practical applications by only incorporating a little training data from the test domain.
V-E More Analysis on MM-Omni3D
Out-of-domain generalization. In this part, we analyze how the training data volume affects the out-of-domain detection accuracy of MM-UniMODE. Specifically, we train MM-UniMODE using only the SUN-RGBD subset in MM-Omni3D, all the indoor subsets in MM-Omni3D, all indoor subsets except SUN-RGBD, and all indoor subsets besides KITTI, respectively. Then, we verify the performances of these models with the SUN-RGBD test set, and the results are reported in Table X. It can be observed that the detector trained based on all the indoor subsets except SUN-RGBD outperforms the detector trained solely with SUN-RGBD, which indicates the promising out-of-domain generalization capability of MM-UniMODE. This observation also suggests that training detectors with more abundant data from more diverse domains generally improves the out-of-domain generalization abilities of the trained detectors.
MM-Omni3D as Pre-training dataset. In this part, we first delve into how pre-training using MM-Omni3D affects the performance on the KITTI test set, and the results are presented in Table XI. The 4 rows of results in this table correspond to the models trained with KITTI, first pre-trained on DD3D [23] and then fine-tuned using KITTI, solely trained with MM-Omni3D, first pre-trained on MM-Omni3D and then fine-tuned using KITTI, respectively. We can observe that MM-Omni3D serves as a more promising pre-training dataset than DD3D, which is a large-scale depth estimation dataset and has been commonly used by previous works. Therefore, besides serving as a multi-modal 3D object detection benchmark, MM-Omni3D also shows promise as a pre-training dataset and has the potential to favor different kinds of 3D perception tasks.
Figure 8: Visualization of the UniMODE and MM-UniMODE detection results in various scenarios. It can be observed that UniMODE achieves promising detection precision and incorporating depth information further boosts the performance.
V-F Visualization
In Fig. 8, we showcase some detection results of UniMODE and MM-UniMODE in various scenarios. It can be observed that both these two detectors can accurately capture 3D object boxes under both complex indoor and outdoor scenarios. UniMODE realizes promising 3D object detection primarily through algorithm design. Then, via integrating the improvement from the data perspective, the detection results of MM-UniMODE are generally more precise due to the adoption of depth sensor information as input.
VI Conclusion
In this work, we have delved into the unified 3D object detection problem in both the monocular and multi-modal settings. In terms of unified monocular 3D object detection, we have proposed a two-stage detection architecture, an uneven BEV grid split strategy, and a domain alignment method to handle the unstable training problem. In addition, a sparse BEV feature projection technique has been developed to alleviate the computing burden. Combining these techniques, a detector named UniMODE has been obtained and it has achieved SOTA performance in the Omni3D benchmark. To explore the unified multi-modal 3D object detection problem, we have built the first corresponding benchmark named MM-Omni3D. Afterwards, we have extended UniMODE to the multi-modal detector MM-UniMODE based on the the proposed module MIC. Extensive experiments have been performed to verify the effectiveness of the various proposed techniques and demonstrated the outstanding performances of UniMODE and MM-UniMODE. In addition, the experimental results also have confirmed MM-Omni3D can serve as a promising pre-training or co-training dataset to improve the generalization capabilities of 3D object detectors.